OpenStudy (anonymous):

What is the magnetic force acting on an electron if its speed is 3.0 × 106 meters/second and the direction is perpendicular to a magnetic field of 0.020 teslas? The value of q = -1.6 × 10-19 coulombs. F = 0 newtons F = -6.0 × 10-15 newtons F = -9.6 × 10-15 newtons F = -3.0 × 10-16 newtons F = -3.2 × 10-21 newtons

OpenStudy (amonoconnor):

Fb= qvBsin(theta) So, q = 1.602 X 10^(-19)C v = 3 X 10^6m/s B = .02T theta= 90 (do you understand where the theta came from?) Fb = qvBsin(theta) = (1.602X10^(-19))(3X10^6)(.02)(sin(90)) = (4.806X10^(-13))(.02)(1) = 4.806X10^(-13)(.02) = 9.612 X 10^(-15) Do you follow all of that ? :)

OpenStudy (amonoconnor):

And I left the negative out of my calculations, but yes, the force is negative because the charge is.

OpenStudy (anonymous):

yeah i got it now, thank you!!

OpenStudy (amonoconnor):

You bet! :)